# Accelerating Vision Diffusion Transformers with Skip Branches

### v2

@@ 7 Evaluation Details
8 Implementation Details
@@
DeepCache [18] is a training-free caching method designed for U-Net-based diffusion models, leveraging the inherent temporal redundancy in sequential denoising steps. It utilizes the skip connections of the U-Net to reuse high-level features while updating low-level features efficiently. Skip-Cache  shares significant similarities with DeepCache but extends the method to DiT models. Specifically, we upgrade traditional DiT models to Skip-DiT  and cache them using Skip-Cache . In the work of DeepCache, two key caching decisions are introduced: (1) N: the number of steps for reusing cached high-level features. Cached features are computed once and reused for the next N-1 steps. (2) The layer at which caching is performed. For instance, caching at the first layer ensures that only the first and last layers of the U-Net are recomputed. In Skip-Cache, we adopt these two caching strategies and additionally account for the timesteps to cache, addressing the greater complexity of DiT models compared to U-Net-based diffusion models. For all tasks except the class-to-image task, caching is performed at the first layer, whereas for the class-to-image task, it is applied at the third layer.
@@
$\Delta$-DiT [6] is a training-free caching method designed for image-generating DiT models. Instead of caching the feature maps directly, it uses the offsets of features as cache objects to preserve input information. This approach is based on the observation that the front blocks of DiT are responsible for generating the image outlines, while the rear blocks focus on finer details. A hyperparameter $b$ is introduced to denote the boundary between the outline and detail generation stages. When $t\leq b$, $\Delta$-Cache is applied to the rear blocks; when $t>b$, it is applied to the front blocks. The number of cached blocks is represented by $N_{c}$.
While this caching method was initially designed for image generation tasks, we extend it to video generation tasks. In video generation, we observe significant degradation in performance when caching the rear blocks, so we restrict caching to the front blocks during the outline generation stage. For Hunyuan-DiT [16], we cache the middle blocks due to the U-shaped transformer architecture. Detailed configurations are provided in Table 9.
@@
Figure 5: Qualitative results of text-to-video generation. We present Skip-Cache , PAB469, and the original model. We apply Skip-DiT on Latte as the backbone model for video generation. The frames are randomly sampled from the generated video.
Figure 6: Qualitative results of class-to-video generation. We present the original video generation model and Skip-Cache with different caching steps $n$. We apply Skip-DiT on Latte as the backbone model for video generation. The frames are randomly sampled from the generated video.
Figure 7: Qualitative results of text-to-image generation. We present Skip-Cache , $\Delta$-DiT, FORA, T-GATE, and the original model. We apply Skip-DiT on Hunyuan-DiT as the backbone model for image generation.
Figure 8: Qualitative results of class-to-image generation. We present the original image generation model and Skip-Cache with different caching steps $n$. We apply Skip-DiT on Hunyuan-DiT as the backbone model for image generation.
@@ 9 Case Study
In Figure 5, we showcase the generated video frames from text prompts with Skip-Cache , PAB, and comparing them to the original model. From generating portraits to scenery, Skip-Cache consistently demonstrates better visual fidelity along with faster generation speeds. Figure 6 presents class-to-video generation examples with Skip-Cache with varying caching steps $\in\{2,4,6\}$. By comparing Skip-Cache to Original model, we see Skip-Cache maintain good generation quality across different caching steps.
@@
Figure 7 compares qualitative results of Skip-Cache compared to other caching-based acceleration methods ($\Delta$-DiT, FORA, T-GATE) on Hunyuan-DiT. In Figure 8, Skip-Cache show distinct edges in higher speedup and similarity to the original generation, while other baselines exist with different degrees of change in details such as color, texture, and posture. Similarly, we present Skip-Cache with varying caching steps in Figure 8, showing that with more steps cached, it still maintains high fidelity to the original generation.

### v4

@@ 9 Evaluation Details
Table 11: Comparison of our caching method with the faster sampler DDIM. Skip-Cache  is evaluated with a 250-steps DDPM and compared to the DDIM sampler under similar throughput. Baseline DiT models (without caching) are highlighted in grey, and the best metrics are indicated in bold. Notably, DDIM outperforms the 250-step DDPM in the UCF101 task for both Latte and Skip-DiT . Skip-Cache  denotes Skip-DiT  with cached inference.
Method UCF101 FFS Sky Taichi
Latte 165.04 23.75 28.88 5.36 49.46 11.51 166.84 11.57
Skip-DiT 173.70 22.95 20.62 4.32 49.22 12.05 163.03 13.55
Skip-DiTn=2 165.60 22.73 23.55 4.49 51.13 12.66 167.54 13.89
DDIM+Skip-DiT 134.22 24.60 37.28 6.48 86.39 13.67 343.97 21.01
DDIM+Latte 146.78 23.06 39.10 6.47 78.38 13.73 321.97 21.86
Skip-DiTn=3 169.37 22.47 26.76 4.75 54.17 13.11 179.43 14.53
DDIM+Skip-DiT 139.52 24.71 39.20 6.49 90.62 13.80 328.47 21.33
[…]
10 Scheduler Selection for text-to-video Tasks
For the text-to-video generation task, all the videos generated for evaluation are sampled with 50 steps DDIM [33], which is the default setting used in Latte. In the class-to-video generation tasks, vanilla Latte uses 250-step DDPM [8] as the default solver for class-to-video tasks, which we adopt for all tasks except UCF101. For UCF101, we employ 50-step DDIM [33], as it outperforms 250-step DDPM on both Latte and Skip-DiT. Table 11 highlights this phenomenon, showing our methods consistently outperform DDPM-250 under comparable throughput, except for UCF101, where DDIM performs better than 250 steps DDPM. In the text-to-image task, we choose 50-step DDIM for sampling, and for the class-to-image task, we choose 250-steps DDPM.
11 Implementation of Other Caching Methods
@@
DeepCache [17] is a training-free caching method designed for U-Net-based diffusion models, leveraging the inherent temporal redundancy in sequential denoising steps. It utilizes the skip connections of the U-Net to reuse high-level features while updating low-level features efficiently. Skip-DiT  shares significant similarities with DeepCache but extends the method to DiT models. Specifically, we upgrade traditional DiT models to Skip-DiT  and cache them using Skip-DiT . In the work of DeepCache, two key caching decisions are introduced: (1) N: the number of steps for reusing cached high-level features. Cached features are computed once and reused for the next N-1 steps. (2) The layer at which caching is performed. For instance, caching at the first layer ensures that only the first and last layers of the U-Net are recomputed. In Skip-DiT, we adopt these two caching strategies and additionally account for the timesteps to cache, addressing the greater complexity of DiT models compared to U-Net-based diffusion models. For all tasks except the class-to-image task, caching is performed at the first layer, whereas for the class-to-image task, it is applied at the third layer.
@@
$\Delta$-DiT [6] is a training-free caching method designed for image-generating DiT models. Instead of caching the feature maps directly, it uses the offsets of features as cache objects to preserve input information. This approach is based on the observation that the front blocks of DiT are responsible for generating the image outlines, while the rear blocks focus on finer details. A hyperparameter $b$ is introduced to denote the boundary between the outline and detail generation stages. When $t\leq b$, $\Delta$-Cache is applied to the rear blocks; when $t>b$, it is applied to the front blocks. The number of cached blocks is represented by $N_{c}$. While this caching method was initially designed for image generation tasks, we extend it to video generation tasks. In video generation, we observe significant degradation in performance when caching the rear blocks, so we restrict caching to the front blocks during the outline generation stage. For Hunyuan-DiT [14], we cache the middle blocks due to the U-shaped transformer architecture. Detailed configurations are provided in Table 12.
@@
Figure 6: Qualitative results of text-to-video generation. We present Skip-DiT , PAB469, and the original model. The frames are randomly sampled from the generated video.
Figure 7: Qualitative results of class-to-video generation. We present the original video generation model and Skip-DiT with different caching steps $n$. The frames are randomly sampled from the generated video.
Figure 8: Qualitative results of text-to-image generation. We present Skip-DiT , $\Delta$-DiT, FORA, T-GATE, and the original model.
Figure 9: Qualitative results of class-to-image generation. We present the original image generation model and Skip-DiT with different caching steps $n$.
@@
AdaCache
AdaCache [11] identifies the feature sample-variance in the DiT, and proposes to predict the feature similarity of the current timestep to decide whether to cache. The codebook for Latte, which records the threshold is not released. We provide a version of codebook which almost reproduce the performance in the paper. The codebook is as follows: $\{0.08:3,0.10:2,0.12:1\}$
TeaCache
Same as AdaCache, TeaCache[16] also discovers the feature instability in DiT. Different from AdaCache, TeaCache finds the relationship between input and output similarity and employ high-order polynomial functions to predict the similarity of the current timestep. We use the official codebase of it and choose the slow-caching strategy to evaluate its upper-bound.
@@ 12 Case Study
In Figure 6, we showcase the generated video frames from text prompts with Skip-DiT , PAB, and comparing them to the original model. From generating portraits to scenery, Skip-DiT  with caching consistently demonstrates better visual fidelity along with faster generation speeds. Figure 7 presents class-to-video generation examples with Skip-DiT with varying caching steps $\in\{2,4,6\}$. By comparing the output of Skip-DiT  with cache to standard output, we see Skip-DiT maintains good generation quality across different caching steps.
@@
Figure 8 compares qualitative results of Skip-DiT compared to other caching-based acceleration methods ($\Delta$-DiT, FORA, T-GATE) on Hunyuan-DiT. In Figure 9, Skip-DiT show distinct edges in higher speedup and similarity to the original generation, while other baselines exist with different degrees of change in details such as color, texture, and posture. Similarly, we present Skip-DiT with varying caching steps in Figure 9, showing that with more steps cached, it still maintains high fidelity to the original generation.